OpenStudy (anonymous):

A non volatile oil has a density of 0.9 g / ml what height of a coloum of oil will represent an atmospheric pressure value of 740 torr (of Hg) density of Hg = 13.6 g/ml and temperature is 298K?

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

I'm sorry that I haven't learnt that... :(

OpenStudy (anonymous):

OK!

OpenStudy (unklerhaukus):

\[1[\text{Torr}]=1[\text{ml}]\cdot[\text{Hg}]=1[\text{cm}^3]\cdot[\text{Hg}]\] \[[\text{Hg}] =13.6 \left[\frac {\text{g}}{{\text{cm}^3}}\right]\]\[[\text {Oil}]=0.9 \left[\frac {\text{g}}{{\text{cm}^3}}\right]\] \[P=740[\text{Torr}]\]\[=740[\text{cm}^3]\cdot1[\text{Hg}]\]\[=h[\text{cm}]\times 1[\text{cm}^2] \cdot[\text {Oil}]\]

OpenStudy (unklerhaukus):

So you are given a pressure, with units that depend on the density of mercury Mercury ;\(\text{Hg} \) can be found in a \(1[\text{cm}^2] \) cylinder i a barometer,